OpenStudy (anonymous):

solve for x, 24/20=x/50

OpenStudy (anonymous):

It's a ratio, just simplify the one fraction to make multiplying easier. 24/20 = x/50 6/5 = x/50 now cross multiply 300 = 5x x = 60

OpenStudy (anonymous):

im still confused..how'd you get it to 60? but how do i know if its a ratio or not?

OpenStudy (anonymous):

Not a problem. Let me rewrite this as nice fractions, so you can visualize it better. You know it's a ratio because you have two fractions set equal to each other, and there is a variable. \[\frac{ 24 }{ 20 } = \frac{ x }{ 50 }\]Because 24 and 20 have a greatest common factor of 4, we can divide each by that to simplify the fraction. \[\frac{ 6 }{ 5 } = \frac{ x }{ 50 }\]Now you could simplify the 6 and 50 if you wanted to as well, because they are diagonal from each other. So let's do that too! \[\frac{ 3 }{ 5 } = \frac{ x }{ 25 }\]Now simply cross-multiply! \[3*25 = 5x \] OR \[75 = 5x\]Now simply divide to finish solving for x. \[x = \frac{ 75 }{ 5 } = 15\]